OpenStudy (anonymous):

Write the equation in standard form. write equation in standard form 4x – 3 = 5y

OpenStudy (anonymous):

4x-5y = 3 Standard form = ax + by= C

OpenStudy (anonymous):

thank you

OpenStudy (nirmalnema):

y=mx+c m = slope. 5y=4x-3 y=4/5x-3/5